### v2

# Bibliography management: biblatex package

Abstract
The perception that the convergence of biological engineering and artificial intelligence (AI) could enable increased biorisk has recently drawn attention to the governance of biotechnology and artificial intelligence. The 2023 Executive Order, Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence, requires an assessment of how artificial intelligence can increase biorisk. Within this perspective, a simplistic framework for evaluating biorisk is presented and discussion is provided on how this framework falls short in achieving actionable outcomes for a biorisk manager. A potential path forward, which builds upon existing risk characterization work, is then suggested with justification to why characterization efforts of AI-enabled tools for engineering biology is needed.
1 Introduction
Biological weapons are a combination of a biological agent and a mechanism to disseminate the agent to a target population. Today, the biological agents are most often considered in the context of Category A agents, as determined by the U.S. Centers for Disease Control and Prevention, and include pathogens that cause Anthrax, Botulism, Plague, Smallpox, Tularemia, and Viral hemorrhagic fevers (e.g., Ebola Virus Disease). The causative agents of these diseases can easily infect large numbers of people because they are person-to-person transmissible or easily disseminated over large areas. These diseases have high mortality rates, would cause public panic, and disrupt day-to-day life.
There is currently an international norm that the use of biological weapons is unacceptable. This norm is codified through the Biological Weapons Convention (BWC), an international treaty. The BWC, which entered into force in 1975, bans Nation States from the development, production, acquisition, transfer, and stockpiling of biological weapons. The BWC is currently signed by 185 Nation States and is complemented by UN Security Resolution 1540 and the Geneva Convention. Resolution 1540 requires all States to prevent the proliferation of biological weapons to non-state groups and the Geneva Convention prohibits use of biological weapons. However, the 1984 Salmonella salad bar attacks by the Rajneesh in the United States, the 1993 anthrax attack attempts by the Aum Shinrikyo cult in Japan, as well as the Amerithrax letter attacks of 2001, serve as reminders that policy measures can only go so far in preventing the misuse of biology. Today, the United States assesses that China and Iran display activities that raise concern regarding biological weapons and that North Korea and Russia maintain active biological weapons programs ([15]).
Around the time the BWC was entering into force, researchers were meeting in Pacific Grove, California to discuss the implications of an important new advance – recombinant DNA technology ([2]). Researchers had recently demonstrated the ability to put DNA from one organism into another, and the scientific community had concerns about the human and environmental health risks in addition to optimism about its potential benefits. Scientists called for a moratorium on such work until the magnitude of risks were assessed and stringent safety guidelines were established. The meeting, now simply known as the Asilomar Conference, named after the conference center, paved the way for responsible governance of recombinant DNA technology. Today, many dual use technologies (those that can be used for both beneficial and malicious purposes) present governance challenges because actions that would mitigate their risks would likely restrict their potential benefits.
The ways in which dual use life science technologies could be misused have been categorized by the late Jonathan B. Tucker and include, 1) “facilitate or accelerate the production of standard biological warfare agents”, 2) “identify or develop novel biological warfare agents”, 3) “result in new knowledge that could be misused to develop a new or enhanced means of harming the human body”, and 4) “lead to harmful applications that undermine international legal norms” ([7]). Since the Asilomar conference in 1975, many technological developments beyond recombinant DNA have demonstrated the potential to be misused and are therefore deemed to have “dual-use potential”. Such technologies include protein engineering, synthesis of viral genomes, gene drives, personal genomics, gene therapy, CRISPR, and immunological modulation. In addition to their benefits, these technologies make it easier for researchers to engineer organisms in ways that make them more suitable to cause harm. A 2004 NASEM report, known as The Fink Report, details seven experimental outcomes that are particularly concerning ([4]). Building upon The Fink Report, the 2018 NASEM report Biodefense in the Age of Synthetic Biology includes a framework for evaluating the relative concern of synthetic biology enabled capabilities ([3]). More recently, the perception that the convergence of biological engineering and artificial intelligence (AI) could enable increased harms has drawn attention to the governance of biotechnology and artificial intelligence ([6, 14]).
2 Biosecurity Risk
President Biden recently signed, Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence. Section 4.4.ii.a of the Executive Order requires the Secretary of Defense, in coordination with many others, to complete a study that, “assesses the ways in which AI can increase biosecurity risks, including risks from generative AI models trained on biological data, and makes recommendations on how to mitigate these risks” ([16]). Within this perspective, a framework is proposed for assessing the ways in which AI can increase biosecurity risks and some observations regarding challenges in implementation of currently proposed recommendations.
Biosecurity risk can be defined as a complex scenario in which a biological organism is intentionally used to cause harm or damage. For the purposes of this work, it is helpful to consider a risk quantification framework as the triplet ($S_{i}$, $P_{i}$, $C_{i}$), where $S_{i}$ is the ith scenario, $P_{i}$ is the probability of that scenario occurring, and $C_{i}$ is the consequence of the ith scenario, where i = 1,2,…N ([1]). To enable comparison, the use of artificial intelligence is assumed to create a parallel set of ai scenarios with probabilities and consequences, denoted ($S_{ai}$, $P_{ai}$, $C_{ai}$), for each of the N scenarios. Each ai scenario can be considered an i scenario that has been modified to include the use of AI. Assuming the biosecurity risk of each of the i scenarios is discrete, AI will increase biosecurity risk when there is an increase in the probability or consequence term of a given ai scenario relative to the corresponding i scenario. Succinctly, AI can increase biosecurity risk through increasing the probability and/or consequence of using biology to cause harm or damage, assuming the paired term (i.e., probability or consequence) does not decrease more than the increase.
3 Probability
For the purposes of this perspective, each scenario is considered to be composed of six steps, denoted $s_{i}^{step}$ or $s_{ai}^{step}$, previously described as the biorisk chain model ([12]). Each of the six steps is described in Table 1 and has an associated probability, $p_{i}^{step}$ or $p_{ai}^{step}$. Use of a model to describe all potential misuse scenarios creates the chance to undervalue or oversimplify components of a misuse scenario. In addition, scenarios of misuse may or may not include all steps in a given biorisk chain model. However, establishing and using biorisk chain models in biorisk analysis enables clear communication around misuse steps, underlying assumptions, and components of the analysis. At present, $S_{i}$ and $S_{ai}$ will have the same steps because AI tools are currently not at a maturity in which they are able to reliably replace a step or obviate the need of a future step.
Table 1: Component steps of the biorisk chain, as defined by Sandberg and Nelson.
Step Description Notation
Insight
Malintent is formed
$s_{i}^{insight}$
Idea
The means or method for causing harm is established; a plan for the following steps is developed
$s_{i}^{idea}$
Sequence
Biological data
$s_{i}^{seq}$
Biosystem
Creation of the physical pathogen
$s_{i}^{bio}$
Testing
Evaluation of the pathogen to cause harm in living things
$s_{i}^{test}$
Dispersal
Release of the tested pathogen to cause harm
$s_{i}^{disp}$
The probability term, $P_{i}$ and $P_{ai}$, is considered to represent the likelihood of successful execution of each step of a sequential plan, spanning ideation through dispersal and regardless of the consequence. For each step, the probability of success can be influenced by many factors including the required resources, skills, and knowledge to perform the step. Probability of success can also be influenced by external policy factors like export controls or DNA synthesis screening regulations. Consequence can be influenced by the steps taken in the plan itself, but also by external factors like preparedness to respond to a plan (e.g., with vaccines or antibiotics). Consequence can also be influenced by human behavior (e.g., mask wearing), policy decisions (e.g., mask mandates), and natural phenomena (e.g., weather). As such, consequence cannot be determined by only the characteristics of the pathogen. Success and consequence can be influenced by the decision to include or exclude steps from the plan.
For example, if a nefarious actor intended to spread Salmonella at a salad bar to prevent community members from voting in a local election, the probability and consequence terms may be altered based on the specific strain of Salmonella used. A more pathogenic strain of Salmonella may be more challenging to acquire, therefore decreasing the probability of successfully executing the overall plan because of a barrier introduced at one step of the plan. At the same time, this decision would increase the consequence term corresponding to severity of disease. But because the disease would be more severe or unusual, it may be more likely to generate a more significant public health response that could decrease the consequence. If the strain of Salmonella were well-characterized, the uncertainties of the consequences would be well understood. However, if the strain were less well-characterized in the environment, perhaps because it was engineered, the uncertainty of the consequence would increase. Given the many component parameters of the probability and consequence terms, they are not independent of each other and will inherently have meaningful associated uncertainty. Therefore, simplifying assumptions to allow only one component of the scenario pair to change may be required to perform a comparison.
This framework can be exercised to identify where AI can increase biosecurity risk. Taking a simplified scenario $S_{i}$ in which a relatively knowledgeable lone actor has an insight to synthesize a pathogen in a lab using nucleic acid material provided by a mail order DNA synthesis company. The specific bottleneck to consider in this comparison is that many mail-order DNA synthesis companies evaluate incoming orders to prevent malicious actors from acquiring potentially harmful material. To overcome this bottleneck, an actor would need to identify a company that would provide them the material – and for this scenario comparison it is assumed that the actor has a limited and fixed amount of time. Within $s_{i}^{idea}$, it can be assumed that the actor uses Google or another search engine. In the comparator scenario step, $s_{ai}^{idea}$, the lone actor uses a non-aligned LLM-based chatbot and Google. If the outcome of this step is assumed to be simply “success” or “fail”, there are then four paired outcomes to consider. To enable a simplistic comparison, “success” is assumed to be reached in the same way (i.e., identifying the same company to order from), to allow for the difference between future $s_{i}$ and $s_{ai}$ to remain negligible.
If each $p_{i}^{idea}$ and $p_{ai}^{idea}$ were quantified, a difference in risks contributed by the scenario pairs of (success — fail) and (fail —success) could be estimated. However, the data to perform this quantification is not available and therefore such assessments rely on expert judgements, although some efforts are underway in an attempt to achieve quantification, including some from model developers ([11]). While these activities are important, they capture a limited number of scenarios and are unlikely to keep pace with advances in AI, such as the growing popularity of retrieval-augmented generation (a technique that supplements an LLM’s internal knowledge with external, predefined sources) and therefore will struggle to produce timely inputs into probability calculations ([9]).
Given the general understanding that AI-powered chatbots make it easier to acquire information that is available online, and that the information needed to overcome the screening bottleneck is available online, it is reasonable to assume $p_{ai}^{idea}$ is greater than $p_{i}^{idea}$ and that AI-powered chatbots therefore increase biosecurity risk in this way. But because the magnitude of the change is not known, it is challenging for a risk manager to evaluate if the increase in risk is meaningful or acceptable.
One approach to determining the acceptability of an increase in risk associated with the impact of AI on a given step in the scenario could be to qualitatively evaluate the relative contribution of each step to the overall scenario and then compare this evaluation with the potential benefits of the AI technology. In the biorisk chain, each step has different limiting factors making the relative contribution of $p_{i}$ to $P_{i}$ variable (Table 2). Given that the magnitude of the risk is not known, experts are often relied upon to conceptualize a meaning of “low/med/high”. This could then allow for an assessment of if the impact of AI on $s_{i}$ results in an $s_{ai}$ with an increased or decreased $p_{ai}$ compared to $p_{i}$, potentially changing the categorical contribution to $S_{i}$.
Table 2: Notional requirement levels for completion of each step within the biorisk chain and associated probability. Note: The entries in this table are judgements made by the author alone. It is not intended to serve as an authoritative source on requirement levels for each step.
Step Time Cost Knowledge Resources Safeguard $p_{i}^{step}$
Insight Low Low Low Low Low High
Ideation Low Low Low Low Low High
Sequence Low Low Low Low Med High
Biosystem High High High High High Low
Testing High High High High High Low
Dispersal Low High High Med High Med
In this biorisk paradigm, concerning impacts of AI would result in changes of $p_{ai}$ from “low” to “medium” or “high” and from “medium” to “high”. Given the notional content of Table 2, impacts of AI on the biosystem, testing, and dispersal steps may be more important to understand and analyze than those at the insight, ideation, or sequence steps.
4 Consequence
The above scenario comparison is somewhat tractable because the consequence term remains fixed. An increase in the consequence term would also result in an increase in biosecurity risk, assuming that the probability term does not decrease commensurately or more. However, such an assumption is likely unwarranted given the current state of AI technologies in the life sciences and biological engineering. A commonly stated concern is that AI tools may be able to design pathogens that are more transmissible and more pathogenic than those found in nature ([13]). These characteristics of a pathogen are often used as a proxy for consequence. However, as alluded to in the salad bar scenario above, the public health response may differ depending on the severity of disease. Therefore, assessments about the consequence of a deliberate biological attack should be accompanied by a significant amount of uncertainty because it is dependent on unknowable human behavior. Furthermore, this scenario assumes that AI-assisted engineering methods will be able to reach a set of viable end-states (e.g., pathogen design) that are not included within the set of viable end-states of traditional biological engineering methods. To date, this assumption remains largely unproven and lacks adequate rationale. Given the ability to explicitly synthesize DNA sequences, there is no reason to think that an AI tool would generate a design that could not theoretically be designed or identified with sufficient rounds of traditional engineering, such as random mutagenesis. For comparison purposes, and until proven otherwise, it is therefore likely better to assume that traditional engineering approaches and AI-assisted engineering approaches are able to reach an outcome of similar consequence.
If, for the sake of comparison, it is assumed that both $S_{i}$ and $S_{ai}$ have the same set of consequences, the impact of purpose-built AI-based engineering tools on the $p_{ai}$ of each $s_{ai}$ compared to the $p_{i}$ of each $s_{i}$ should be explored, similar to the evaluation of the impact of the chatbots in the prior example. Consider a scenario in which a nefarious actor wants to engineer a pathogen using existing bioengineering methods and the paired scenario containing the use of AI. In this scenario pair, it is assumed that the use of genetic engineering within the biosystem step that requires a traditional synthetic biology process of design – build – test – learn, with the end goal of optimizing the characteristic properties of a pathogen (not the pathogen’s effects on its host) is used ([8]). There is a biomolecule, or “candidate sequence”, that has desirable properties that could be optimized. In $s_{i}^{seq}$ it is assumed the use of random mutagenesis or equivalent method to identify the variants of the candidate sequence to build and test, whereas in $s_{ai}^{seq}$ an AI system generates the recommended sequences to build and test. For the purposes of this evaluation, it is assumed that the requirements to perform mutagenesis and AI design are equal(although, in practice this is likely not true). In $S_{i}$ and $S_{ai}$, $s^{build}$ and $s^{test}$ are identical because they are independent of the AI tool. It is also assumed the $s^{learn}$ is equal between $S_{i}$ and $S_{ai}$, although the specific process of learning from data would proceed differently in each. Therefore, the contribution of AI comes from potentially decreasing the overall number of design-build-test-learn cycles which would reduce the time, cost, and resources required to achieve the desired end goal. To support risk analysis, an understanding of the difference in number of design-build-test-learn cycles between a typical $S_{i}$ and $S_{ai}$ (as a means to characterize a difference in probability of success, with decreasing probability associated with increasing numbers of cycles) is required.
Currently, there is not enough information in the published literature to support a reasonable, generalized characterization of this difference. Often, when researchers are developing AI systems and tools for engineering biology, they will perform the same number of design-build-test-learn cycles to demonstrate that the AI method is improved over the traditional engineering method. For example, a previously reported machine learning method provided a nearly 29-fold improvement compared to a traditional method ([10]). However, it took the team 12+ months to develop the machine learning method that provided the designs. If accounting for the time (and cost, knowledge, resources, etc.) a more apt comparison may be to the number of traditional engineering cycles that could have occurred over that 12+ month development period. Additionally, the authors tested their methods in two very similar engineering tasks and notably, the method that resulted in our best performance in the first scenario gave the worst performance in the second scenario without improving over the traditional engineering method. And lastly, the authors did not perform additional traditional engineering cycles to determine the number of cycles it would take to reach performance of the best AI outcome. To fully understand the impact of including AI methods and tools in biological engineering processes in a manner that supports risk assessment, dedicated efforts to characterize the performance of AI methods across diverse tasks is needed.
Efforts to perform characterization of the ability of AI tools to assist in the engineering pathogens are likely to be controversial. Given that the output of an AI-enabled bioengineering tool is a prediction, validation of the prediction in a biological laboratory is necessary to determine how often the predictions are correct and if there are circumstances or conditions, such as amount or type of training data, that influence the rate of success. Unlike chemical threats, biological threats can transmit from person-to-person and self-amplify. As such, a small mistake in a laboratory could result in large-scale consequence. With this in mind, there should be limits to the extent of characterization of the ability of AI tools to assist in the engineering of pathogens for the sole purpose of informing risk analysis. Creating and evaluating biological agents that are predicted to be highly pathogenic should not be done for the sole purpose of characterizing the predictive capabilities of these tools. This should be the norm, and again, it should not be expected that a quantitative understanding of the impact of AI on biorisk to be achievable.
Proxy evaluation metrics for the ability of AI tools to deliberately engineer pathogens could offer a possible path forward. However, it is unlikely the biosecurity and associated policy communities will be able to agree on sufficient proxy metrics given longstanding disagreements about the ability to “reasonably anticipate” the outcome of scientific experiments ([5]). Therefore, it may be better to consider the capability of an AI tool as a clearly stated assumption in risk analysis conversations or to seek a relative understanding of the level of concern associated with capabilities and scenarios in which biology is used to cause harm. However, the notable number of assumptions required to use the triplet framework described herein may limit its utility in quantitative assessments.
5 A Path Forward
The NASEM Report Biodefense in the Age of Synthetic Biology offers a viable starting point for performing a qualitative risk assessment as required by the recent Executive Order. The report provides relative levels of concern for a set of synthetic biology enabled capabilities; these capabilities will be impacted by artificial intelligence tools, some more so than others. Of particular relevance today are the capabilities of making existing bacteria and viruses more dangerous. The NASEM report includes a list of nine viral traits and eight bacterial traits that could be engineered to result in a more dangerous pathogen. With this in mind, the following steps could be taken to update the evaluations within the NASEM report:
  1. 1.
    Identify organisms of interest. Organisms may biologically differ in the way in which their traits of interest are determined. As such, the way these traits are engineered may be different and thus the AI-enabled tools used to support engineering may be different. This list of organisms could include pathogens with pandemic potential, those that are likely candidates for misuse (e.g., Category A), and those that are often engineered in beneficial biotechnological applications because they likely have existing AI-enabled engineering tools.
  2. 2.
    Identify existing AI-enabled tools that are amenable to engineering the traits of the organisms of interest through literature and web search. Some tools may be purpose-built for engineering whereas other tools may be designed to characterize organisms and in some scenarios could likely be repurposed or modified to engineer the characterized trait.
  3. 3.
    When AI-enabled tools do exist, systematically and qualitatively evaluate their impact on the four categories that, collectively, contribute to the level of concern that should be attributed to a particular scenario. The four categories are Usability of the Technology, Usability as a Weapon, Requirements of Actors, and Potential for Mitigation, with each having multiple sub components. Risk assessors should seek to understand the common uses of the AI tools and how the tools impact the sub components, such as ease of use or required expertise.
  4. 4.
    Given the findings of Step 3, review and update, if necessary, the overall relative level of concern for each capability of concern in the NASEM report.
Similar processes could also be developed and executed for assessing the impact of AI on the capabilities that do not include the engineering of pathogens, such as those that modify the human microbiome, immune system or genome. In aggregate, the outcome of this potential work would be a relative comparison of levels of concern about the capabilities. It would not necessarily allow for a conclusion about overall changes in the absolute level of biorisk, but rather an understanding of how AI impacts potentially concerning scenarios in a relative manner. Even without an absolute quantitative understanding of biorisk, a relative helps risk managers devote the appropriate attention and resources to mitigate biorisk.
As AI-enabled biological engineering tools continue to mature, it will be increasingly important to understand the ways in which they impact the deliberate biological threat landscape. Such risk analysis efforts should be mindful to ensure that the results are useful to risk managers. Given the complexity of the misuse of biology, risk managers should expect and be supportive of relative, qualitative risk analysis that is informed by a technical understanding of the ways in which AI tools are used within biology and the life sciences. This technical understanding can be furthered by dedicated characterization efforts of AI-enabled tools outside of pathogen engineering context. This work is noticeably absent at present.
6 Acknowledgements
The author thanks Dr. Gerald Epstein and Dr. Gigi Kwik Gronvall for their constructive comments and suggestions on the manuscript. The author is funded by Open Philanthropy.